40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Nov 4 00:10:43 2020, gautam, Update, General, Vent activities today IMG_0184.jpgIMG_0185.jpgIMG_0187.jpgIMG_0207.jpegIMG_0210.jpg
    Reply  Thu Nov 5 22:27:34 2020, gautam, Update, General, Vent activities today 
       Reply  Fri Nov 6 11:36:37 2020, gautam, Update, General, ETMY suspension eigenmodes ETMY_pkFitNaive.pdfETMY.tar.bz2
          Reply  Tue Nov 10 15:13:41 2020, rana, Update, General, ETMY suspension eigenmodes 
             Reply  Thu Nov 12 14:26:35 2020, gautam, Update, General, ETMY suspension eigenmodes ETMY.tar.bz2
       Reply  Sun Nov 8 20:22:35 2020, gautam, Update, General, Vent activities today IFOalignment.png
          Reply  Sun Nov 8 23:24:29 2020, Koji, Update, General, Vent activities today 
          Reply  Mon Nov 9 17:40:17 2020, gautam, Update, General, Pumpdown IMG_8937.jpgpumpdown.png
             Reply  Tue Nov 10 11:31:13 2020, Koji, Update, General, Pumpdown Screen_Shot_2020-11-10_at_11.30.21.png
Message ID: 15664     Entry time: Sun Nov 8 20:22:35 2020     In reply to: 15660     Reply to this: 15665   15666
Author: gautam 
Type: Update 
Category: General 
Subject: Vent activities today 

Basic IFO alignment checks were done.

  1. IMC could be locked - I tweaked the cavity alignment a little to maximize the MC transmission.
  2. Y arm can be made resonant for a TEM00 mode of the main beam. I can't run ASS successfully in this low power config. I can see that the axis isn't great, the spot is visually off-center on ITMY, but we should have plenty of actuator range to correct for this with TT1/TT2.
  3. X arm shows IR mode flashes in the TRX QPD. The green beam can be made resonant for a TEM00 mode, but that alignment doesn't yield the largest IR resonant peaks in TRX. I suspect it is due to the mis-alignment of the beam axis.
  4. AS beam was aligned onto the CCD. I could see clean Michelson fringes by tweaking the BS alignment. On the AP table, I noticed that the beam on the first steering mirror after the AS beam exits the vacuum is a little high. We can easily resolve this by tweaking OM6 pitch a bit, but even if we don't I don't see any major issues as there is plenty of clearance w.r.t. the viewport when the beam exits the vacuum.
  5. With the PRM aligned, I can see the REFL beam on the CCD, and it doesn't look clipped.
  6. I didn't bother to align the green beams to the arm cavities or re-center the Oplevs - is this necessary? It is a step in the pre-close up checklist, so maybe we should do it... The green transmission does reach the PSL table...

Tomorrow, we should do some visual checks of the chambers / EQ stops on ETMY etc but I don't see any major problems at the moment...

Quote:
 

Barring any catastrophic failures and provided all required personnel are available, we will do the final pre-close-up checks, put the heavy doors back on, and pump down starting 10 am Monday, 9 Nov 2020.

Attachment 1: IFOalignment.png  316 kB  | Hide | Hide all
IFOalignment.png
ELOG V3.1.3-